OpenStudy (anonymous):

can someone please answer. Solve the equation for x and round your answer to the nearest hundreth. 5 = ln(x + 1)

OpenStudy (anonymous):

x+1=e^5 => x=e^5 -1

OpenStudy (anonymous):

yup, take the base of e or whatever u call it to cancel out the LN.